OpenStudy (anonymous):

Identify the 35th term of an arithmetic sequence where a1 = –7 and a18 = 95.

sam (.sam.):

First find 'd', from a18, a18=a1+(n-1)d a18=(-7)+(17)d=95 d=6 a35=a1+(n-1)d a35=(-7)+(34)6=?